OpenStudy (anonymous):

Here is one that is a little different i am almost getting this and appreciate any help multiply 9xy(x+y)

jimthompson5910 (jim_thompson5910):

9xy times x = ???

OpenStudy (phi):

"distribute" by multiplying 9xy times each term inside the parens 3*( 1+2)= 3*1 + 3*2 = 3+6=9 or 3* 3= 9 same idea, but with variables